OpenStudy (anonymous):

5 + 2{x - 4[3x + 7(2 - x)]} 34x - 107? please dont make me show work if its correct

OpenStudy (bahrom7893):

divided by?

OpenStudy (mathmate):

34x - 107 is correct if the problem is to simplify the expression.

OpenStudy (anonymous):

yes tank you

OpenStudy (bahrom7893):

oh you already had the answer.

OpenStudy (mathstudent55):

Instead of making you do the work, I did it myself. \(5 + 2{x - 4[3x + 7(2 - x)]}\) \(=5 + 2{x - 4[3x + 14 - 7x]}\) \(=5 + 2{x - 4[-4x + 14]}\) \(=5 + 2{x + 16x - 56}\) \(=5 + 2{17x - 56}\) \(=5 + 34x - 112\) \(=34x - 107\) You are absolutely correct. Good job!